OpenStudy (anonymous):

Train A leaves the station at 30mph. Three hours later train B leaves going 40 mph. How far out are the trains when they meet?

OpenStudy (anonymous):

distance A=30t distance B=40(t-3)

OpenStudy (anonymous):

30t=40(t-3) find t

OpenStudy (anonymous):

t=12 hours distane A=30(12)=360 miles

OpenStudy (anonymous):

@Luis_Rivera i think instead of 50 its 40